OpenStudy (anonymous):

m and n are positive integers and leave remainder 2 and 3 respectively when divided by six m>n,,,,,so what would be remainder when m+n is divided by 6?

OpenStudy (anonymous):

m/6 = a+2 n/6 = b+3 m+n/6 = m/6+n/6 = a+b+5 so the remainder is 5.

OpenStudy (anonymous):

thank you !!!! know i got it ..:)))

OpenStudy (anonymous):

Oh God, someone please introduce this poor child to some basic modular arithmetic.

OpenStudy (anonymous):

i am just 4th grader

OpenStudy (anonymous):

here the modulus is 6. I don't know if they teach that in 4th grade.